OpenStudy (anonymous):

4/7+1 ________ 2/7+1 simplify complex fraction

OpenStudy (zehanz):

write 1 as 7/7. You then can add 4/7 and 7/7, so you get one fraction in the numerator. Do the same with 2/7 and 7/7 in the denominator. Then you are ready for the next step...

OpenStudy (anonymous):

ok i got 11/7 and 9/7 what do i do next

OpenStudy (zehanz):

You have: \(\dfrac{\frac{11}{7}}{\frac{9}{7}}\). You can do two things: 1. Multiply 11/7 and 9/7 with 7, and see what happens. 2.Instead of dividing by 9/7, you can also multiply with 7/9: \(\dfrac{11}{7} \cdot \dfrac{7}{9}=...\)

OpenStudy (anonymous):

would it be 11/9 ?

OpenStudy (zehanz):

Yes, it would!

OpenStudy (anonymous):

thanks so much!

OpenStudy (zehanz):

YW!